OpenStudy (anonymous):

Find the length of the segment with the given endpoints. (8,1)and (2,9)

OpenStudy (m):

use the distance formula

OpenStudy (anonymous):

pythagoras gives \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

\[\sqrt{6^2+8^2}=\sqrt{100}=10\]

OpenStudy (anonymous):

Thanks!